Question 2 pts You want to take a cruise vacation 4 years from now when you complete graduate school. You estimate that you will need $5,000 when you complete graduate school to pay for the cruise. If you can earn 9.8% interest on your deposits, how much would you have to deposit today to be able to pay for the cruise when you graduate? Round your answer to 2 decimals. For example, $100.12.D Question 2 2 pts You will deposit 300 per year into an account that earns 7.9% per year. How much will you have in the account 6 years from now? Round your answer to 2 decimals, for example 100.21. Question 3 2 pts You will receive $300 per year for the next 5 years. What is the PV of the payments if the interest rate is 5.7%? Roiund your answer to 2 decimals, for example 100.12
Answer #1

Q1. Amount required to deposit Now to pay for cruise four years later

Amount to deposit = Amount required / (1 + Required Rate)^4

Amount to deposit = 5000 / (1 + 0.098)^4

Amount to deposit = 5000 / 1.45348

Amount to deposit = $3440.02

Q2. Account Balance after 6 Years = Amount deposited * (1 + r)^n

Account Balance after 6 Years = 300 * (1 + 0.079)^6

Account Balance after 6 Years = 300 * 1.5781

Account Balance after 6 Years = $473.42

Q3. Present Value = Receipt per year * PVAF ( 0.057, 5)

Present Value = 300 * 4.24697

Present Value = $1274.09